AFWEX Status, 06 December 2000

Today was a definite fly day for AeroSapient activities on the DC-8. This consisted of a 1800 CST take-off followed by 3 hours of maneuvers slightly northwest of the CART site (see DC-8 nav plots for details) at 39 kft. The weather cooperated and we had mostly clear skies into the evening with thin cirrus at ~11.5 km and so the flight was extended another ~5 hours for AFWEX maneuvers. This consisted of a combination of spiral ascents and descents, level legs both with-wind and cross-wind, and a wide spiral down (allowing for the LASE to operate). The Proteus performed a ~6 hour coordinated flight over the CART site consisting of slow ascents and descents and a mapping pattern at ~54 kft. The Proteus was targeting clear skies and cirrus on this flight, and would still like to obtain a dataset with thicker cirrus. All instruments (aircraft and CART site) performed well. This appears to be another outstanding case for the AFWEX objectives.

At today's 3pm science meeting, preliminary results from the 12/04 CST "severe clear" flight were shown. Rich Ferrare showed a comparison of colocated LASE, CART Raman Lidar, Cryo in-situ, and WFF chilled mirror sonde profiles versus a Vaisala sonde profile; this Vaisala sonde showed a dry bias with respect the other profiles above 6 km. Rich commented that when all of the datasets are compiled, an additional ~5 profiles can be added to this plot. Dave Whiteman showed comparisons of GSFC Raman, CART Raman, and Vaisala sonde profiles, with the the two lidar systems agreeing well in all cases, and with a slight moist bias with respect to the Vaisala sonde at ~10 km. Dave Tobin showed good agreement between a temperature profile derived from SHIS radiance spectra during a spiral down maneuver to those from coincident radiosondes.

Skies became mostly clear at mid-day and into the evening. Beginning ~1800 CST, some thin cirrus moved south over northern Oklahoma. Patchy thin cirrus at 11-12 km was seen for the duration of the flight, but is not expected to impact the water vapor measurements and comparisons.

Instrument Status/Comments
CART site:
   CART Raman Lidar Normal operations. Had a power failure earlier in the day and the system was back on line at ~1330 CST.
   GSFC Raman Lidar Normal operations from 2030 to 0200 CST.
   MPI DIAL Operated from ~1830 to 0200 CST. Excellent data set.
   WFF Chilled Mirror Sondes Launches at 2015, 2315, 0115 CST.
   Other Vaisala sondes at 1730, 2030, 2330 ... CST.
   LASE Normal operations.
   DLH Normal operations.
   Cryo frost-point/chilled-mirror Normal operations.
   SHIS Normal operations.
   COAST Normal operations. No COAST specific flight patterns.
   AeroSapient See Recap above.
   NAST-I Normal operations.
   NAST-M Normal operations.
   FIRSC Normal operations.

Dave Tobin, University of Wisconsin